# The Stellar Content of the Young Open Cluster Berkeley 50 (IC 1310)

### v1

@@ 2.1 Photometry
Images were taken on the LDT using the Large Monolithic Imager (LMI). The imager consists of a single 92.2-mm by 92.4-mm e2v CCD231-C6 chip with a multi-layer anti-reflective coating. It is operated at -120 C using a Stirling closed-cycle cooler. The 125-mm Bonn shutter provides better than 1% uniformity for exposures as short as 0.1 s, allowing broad-band photometry of even 10th magnitude stars in sub-arcsecond seeing. Additional details can be found in the instrument manual.222http://www2.lowell.edu/users/massey/LMIdoc.pdf The CCD was binned 22, resulting in a scale of 024 per binned pixel. After trimming, the resulting field-of-view was 123123. The chip was read out through a single amplifier, with a gain of 2.89 electrons per analog-to-digital unit (ADU). The device has a read-noise of 6.0 electrons. The instrument is linear to better than 1% up to a data value of 60,000 ADU (175,000 electrons) per unbinned pixel.
@@ 2.2 Spectroscopy
All the spectroscopy was carried out on the LDT using the DeVeny optical spectrograph. The detector is a 2048512 e2v CCD42-10 deep depletion device with 13.5 m pixels. It is operated at a gain of 1.5 electrons ADU, and has a full-well of about 100,000 electrons. A Stirling closed-cycle cooler chills the chip to C. A slit viewing camera is used to position stars on the slit. The spatial scale on the LDT is 034 pixel.
A 600 line mm grating blazed at 4900 Å was used with a grating tilt of 2695, which provided wavelength coverage from 3700-5950 Å and a dispersion of 1.12 Å pixel. The slit was opened to a width of 10, which provided a spectral resolution of 2.5 Å. Exposure times of 300-1800 s were used, depending upon the brightness of the object; longer exposures were broken up into three pieces in order to more easily remove radiation events (“cosmic rays”). Spectrophotometric standards were observed at the start of each night to allow flux calibration. All observations were made by rotating the spectrograph so the slit was oriented to the parallactic angle. Wavelength calibration was provided by an exposure of Hg-Cd-Ar comparison lamps at the start of the night. The spectrograph suffers from some flexure, but, unfortunately, the procedure for taking comparison exposures is time-consuming, making it impractical to take comparison spectra for each object. Thus, our data are not good enough for accurate radial velocity information. Dome flats and bias frames were taken in the afternoon.
@@ 3 Membership
In order to establish the likely proper motions of members, we began by considering only those stars within a 5′ radius of the RSG. The median proper motion of this sample is 6.11 mas yr. Within this sample, the median proper motion in right ascension and declination to be = -3.32 mas yr and = -5.13 mas yr respectively. (These values agree well the statistical determination of the proper motions by Hunt & Reffert 2023 who found = -3.37 mas yr and = -5.13 mas yr.) Based on the distribution of values around these values, we chose to restrict the sample to those stars whose proper motions were within 3 of these median values, based on the tabulated Gaia uncertainties. Implementing these proper motion restrictions eliminated 86% of the stars, and reduced the sample of potential members to 163.

### v2

@@ 2.1 Photometry
Images were taken on the LDT using the Large Monolithic Imager (LMI). The imager consists of a single 92.2-mm by 92.4-mm e2v CCD231-C6 chip with a multi-layer anti-reflective coating. It is operated at -120∘ C using a Stirling closed-cycle cooler. The 125-mm Bonn shutter provides better than 1% uniformity for exposures as short as 0.1 s, allowing broad-band photometry of even 10th magnitude stars in sub-arcsecond seeing. Additional details can be found in the instrument manual.222http://www2.lowell.edu/users/massey/LMIdoc.pdf The CCD was binned 22, resulting in a scale of 024 per binned pixel. After trimming, the resulting field-of-view was 123123. The chip was read out through a single amplifier, with a gain of 2.89 electrons per analog-to-digital unit (ADU). The device has a read-noise of 6.0 electrons. The instrument is linear to better than 1% up to a data value of 60,000 ADU (175,000 electrons) per unbinned pixel.
@@ 2.2 Spectroscopy
All the spectroscopy was carried out on the LDT using the DeVeny optical spectrograph. The detector is a 2048512 e2v CCD42-10 deep depletion device with 13.5 m pixels. It is operated at a gain of 1.5 electrons ADU-1, and has a full-well of about 100,000 electrons. A Stirling closed-cycle cooler chills the chip to C. A slit viewing camera is used to position stars on the slit. The spatial scale on the LDT is 034 pixel-1.
A 600 line mm-1 grating blazed at 4900 Å was used with a grating tilt of 2695, which provided wavelength coverage from 3700-5950 Å and a dispersion of 1.12 Å pixel-1. The slit was opened to a width of 10, which provided a spectral resolution of 2.5 Å. Exposure times of 300-1800 s were used, depending upon the brightness of the object; longer exposures were broken up into three pieces in order to more easily remove radiation events (“cosmic rays”). Spectrophotometric standards were observed at the start of each night to allow flux calibration. All observations were made by rotating the spectrograph so the slit was oriented to the parallactic angle. Wavelength calibration was provided by an exposure of Hg-Cd-Ar comparison lamps at the start of the night. The spectrograph suffers from some flexure, but, unfortunately, the procedure for taking comparison exposures is time-consuming, making it impractical to take comparison spectra for each object. Thus, our data are not good enough for accurate radial velocity information. Dome flats and bias frames were taken in the afternoon.
@@ 3 Membership
In order to establish the likely proper motions of members, we began by considering only those stars within a 5′ radius of the RSG. The median proper motion of this sample is 6.11 mas yr-1. Within this sample, the median proper motion in right ascension and declination to be = -3.32 mas yr-1 and = -5.13 mas yr-1 respectively. (These values agree well the statistical determination of the proper motions by Hunt & Reffert 2023 who found = -3.37 mas yr-1 and = -5.13 mas yr-1.) Based on the distribution of values around these values, we chose to restrict the sample to those stars whose proper motions were within 3 of these median values, based on the tabulated Gaia uncertainties. Implementing these proper motion restrictions eliminated 86% of the stars, and reduced the sample of potential members to 163.